# Solving the Gross-Pitaevskii Equation with Quantic Tensor Trains: Ground States and Nonlinear Dynamics

### v2

@@ Appendix D Convergence behavior
Here we discuss more details about the convergence properties in the QTT methods. In the variational method, we minimize the energy $E[\psi]$ rather than the chemical potential $\mu$. Therefore, $\mu$ can be lower than the exact value during the optimizations. In Fig. 12(a), we show the convergence of the energy for the system of BEC in a harmonic trap. It can be seen that the energy decreases almost monotonically and is always higher than the exact value during the optimization steps. In some steps, the energy jumps up and then decreases again. This is because the gradient in our method is computed approximately and could sometimes allow an increase in energy. However, this does not slow down the overall convergence in the algorithm. In Fig. 12(b), we show the energy slopes along the searching directions as functions of optimization steps. Similar to the energies, the slopes decrease rapidly to almost zero with optimization steps, showing convergence of the simulations.

### v3

@@ Appendix D Convergence behavior
Figure 13: Absolute errors in the chemical potential (a) and total energy (b) as functions of the number of qubits for each dimension for a system of a BEC in a harmonic trap.
@@
In this section, we discuss the convergence properties of QTT methods in more detail. We demonstrate on a system of a BEC in a harmonic trap (see Fig. 3(a)). We first show the absolute errors of chemical potential $\mu$ and energy $E$ as functions of the number of qubits for each dimension in Fig. 13. (The number of grid points grows exponentially with the number of qubits.) As expected, the errors decay exponentially with the number of qubits, showing the efficiency of the QTT approach.
We then discuss the convergence properties in the variational method. In the variational method, we minimize the energy $E[\psi]$ rather than the chemical potential $\mu$. Therefore, $\mu$ can be lower than the exact value during the optimizations. In Fig. 14(a), we show the convergence of the energy for the system of BEC in a harmonic trap. It can be seen that the energy decreases almost monotonically and is always higher than the exact value during the optimization steps. In some steps, the energy jumps up and then decreases again. This is because the gradient in our method is computed approximately and could sometimes allow an increase in energy. However, this does not slow down the overall convergence in the algorithm. In Fig. 14(b), we show the energy slopes along the searching directions as functions of optimization steps. Similar to the energies, the slopes decrease rapidly to almost zero with optimization steps, showing convergence of the simulations.